# Quantifier Elimination and Invariant Theory: Applications to Quaternions, Octonions, and Other Algebras

### v1

@@ 1.1 Introduction
In contrast, much less is known about quantifier elimination for noncommutative algebras. Even for natural and well-studied structures such as the classical quaternions $\mathbb{H}(k)$ or octonions $\mathbb{O}(k)$, for $k$ real closed, the model-theoretic behavior is considerably more subtle. In particular, it was recently shown in [savi] that their first-order theories in the pure ring language do not admit quantifier elimination, and the problem of identifying suitable language extensions was left open.
@@
As applications, we prove quantifier elimination for the classical quaternions and octonions over real closed fields in a natural language expansion by conjugation $\overline{}$ and order $\leq$. More generally, we obtain criteria for finite-dimensional algebras $\mathcal{A}$ over real closed and algebraically closed fields, connecting properties of their invariant algebras $k[\mathcal{A}^{m}]^{G}$ for $G<\text{Aut}(\mathcal{A})$, to possible candidates of language extensions of $\mathcal{F}$, such that quantifier elimination holds.
@@ 1.4 Linear Algebraic Groups and Invariant Theory
Let $k$ be an algebraically closed field of characteristic 0 and $G<\mathrm{GL}_{n}(k)$. We say that $G$ is a linear algebraic group (over $k$) if it is Zariski closed in $\mathrm{GL}_{n}(k)$, i.e. if there exists an ideal $I$ of polynomial functions in the matrix entries such that
@@
The next theorem, in the form we use it, reduces to an application of the Stone–Weierstrass theorem.
@@ 2 Results
The $(\mathcal{F},\leq,\overline{\ })$-structure $\mathbb{H}(k)$ admits quantifier elimination.
@@
We apply Theorem˜2.1 to the $(\mathcal{F},\leq,\overline{\ })$-structure $\mathbb{H}(k)$ and the $(\mathcal{F},\leq)$-structure $k$, which admits quantifier elimination.
@@
each coordinate of $\iota^{(m)}(f_{m}(y_{1},...,y_{m}))$ is a simple $(\mathcal{F},\overline{\ })$-term. Therefore
@@
is a quantifier-free $(\mathcal{F},\leq,\overline{\ })$-formula defining $f_{m}^{-1}(D)$.
@@
$\mathrm{Aut}(\mathbb{H}(k))_{(\mathcal{F},\leq,\overline{\ })}=\mathrm{Aut}(\mathbb{H}(k))_{\mathcal{F}}=SO_{3}(k).$
@@
The $(\mathcal{F},\leq,\overline{\ })$-structure $\mathbb{O}(k)$ admits quantifier elimination.
@@
For both quaternions and octonions the symbol $\overline{\ }$ was needed to be able to express the invariant in our extended language. Note that $\overline{\ }$, $\operatorname{Re}$, $\operatorname{Im}$, and $N$ are interdefinable, so the precise choice of the additional symbol can be handled flexibly. On the other hand it was essential for our proof to extend $\mathcal{F}$ by $\leq$ and interpret it as the partial order on the center. This had to be done to ensure condition (ii) of Theorem˜2.1 and indeed, the symbol $\overline{}$ allows for defining a copy of $k$ in the $(\mathcal{F},\overline{})$-structure $\mathbb{H}(k)$ or $\mathbb{O}(k)$ without quantifiers. Thus formulas like
@@
Then, since $G$ is a compact subgroup of $\mathrm{Aut}(\mathcal{A})<\mathrm{GL}_{n}(k)$ it is a Lie group and acts diagonally on $\mathcal{A}^{m}$. Thus by Theorem˜1.5 $k[\mathcal{A}^{m}]^{G}$ separates the orbits of the diagonal action $G\curvearrowright\mathcal{A}^{m}$. Since $(s_{i})$ generate $k[\mathcal{A}^{m}]^{G}$ we get the following relation:
@@
The conditions for the language extension $E$ Theorem˜2.2 seem somewhat far fetched. However, note that so called first fundamental theorems (see [sym]) give explicit descriptions of generators for many invariant algebras of $k$-algebras, and their corresponding automorphism groups (or subgroups thereof) by means of tensor forms. These are usually stated for the algebraically closed case, however, it should be possible to often use Zariski denseness arguments, as we did in the proof of Lemma˜2.2, to infer back to the real closed case.
What about algebras $\mathcal{A}$ over an algebraically closed field (in the following always denoted $k$)? Compactness of the acting group was essential in the real closed case. Comparing with Theorem˜1.2, we see that in general $\mathrm{Aut}(\mathcal{A})$ (or subgroups thereof) separate only orbit closures, which need not separate $\emptyset$-types. Unfortunately compactness for linear algebraic groups $G$ over $k$, only holds in the case of $G$ finite, restricting us quite a bit.
@@
First note that by Theorem˜1.3 all $S_{m}$ are finite and so define
@@
Since $k$ is algebraically closed, it admits quantifier elimination. Conditions (i) and (ii) of Theorem˜2.1 are verified exactly as in Theorem˜2.2.
For (iii), note that since a finite group is discrete it is by definition reductive and has closed orbits. Hence by Theorem˜1.2, $k[\mathcal{A}^{m}]^{G}$ separates $G$-orbits. As before,
@@
For an associative $k$-algebra $\mathcal{A}$, by an application of Wedderburn’s Theorem $\text{Aut}(\mathcal{A})$ is only finite in the case $\mathcal{A}\cong\mathbb{C}^{n}$. However finite subgroups of $\text{Aut}(\mathcal{A})$ exist and their invariant algebras, can be studied ([comp, fin]). For non-associative algebras, $\text{Aut}(\mathcal{A})$ can even itself be finite (and non-trivial). For example in [low], all semisimple, two dimensional $k$-algebras have been classified and their corresponding automorphism groups (some of which are finite) and invariant algebras calculated explicitly.
We finish with a statement using compact $\mathbb{R}$-rational points of linear algebraic groups without the necessity to extend the language by $\leq$. For simplicity we write it this time for $\mathbb{R}$ and $\mathbb{C}$. Generalization should again be possible, by translating everything into notions that can be transferred as first order logic statement.
Theorem 2.4.
Let $\mathcal{A}$ be a finite-dimensional $\mathbb{C}$-algebra. Assume $G<\mathrm{Aut}(\mathcal{A})_{\mathcal{F}}$ is a connected reductive linear algebraic group defined over $\mathbb{R}$ and let the $\mathbb{R}$-rational points $G_{\mathbb{R}}$ be compact. Let $S_{m}$ be a minimal generating set of $\mathbb{C}[\mathcal{A}^{m}]^{G}$ for each $m\in\mathbb{N}$, and let $(\mathcal{F},E)$ be an extended language such that the interpretation of each $e\in E$ in $\mathcal{A}$ lies in $\iota(\mathbb{C}[\mathcal{A}^{\text{arity}(e)}]^{G})$ and for all $s\in S_{m}$, $\iota(s)$ corresponds to evaluation of a $(\mathcal{F},E)$-term. Then the $(\mathcal{F},E)$-structure $\mathcal{A}$ admits quantifier elimination.
Proof.
By reductiveness of $G$ and Theorem˜1.3, each $S_{m}$ is finite, so define
$f_{m}:\mathcal{A}^{m}\to\mathbb{C}^{|S_{m}|},\qquad f_{m}(\underline{a})_{i}=s_{i}(\underline{a}).$
We want apply Theorem˜2.1 from the $(\mathcal{F},E)$ structure $\mathcal{A}$ to the $\mathcal{F}$ structure $\mathbb{C}$.
Condition (i) and (ii) can be shown exactly as in the previous proofs.
The only non trivial part again is (iii). By definition on $G$ Theorem˜1.2 implies:
$f_{m}(\underline{a})=f_{m}(\underline{b})\Rightarrow i(\underline{a})=i(\underline{b})\forall i\in\mathbb{C}[\mathcal{A}^{m}]^{G}\Leftrightarrow\overline{\mathrm{Orb}(\underline{a},G)}\cap\overline{\mathrm{Orb}(\underline{b},G)}\neq\emptyset$
So take a sequence $(g_{n})_{n\in\mathbb{N}}\subset G$, such that $g_{n}\cdot\underline{a}\to\underline{c}\in\overline{\mathrm{Orb}(\underline{b},G)}$. As $G$ is defined over $\mathbb{R}$ by the Hilbert-Mumford stability criterion (see [GIT] and [KN]) if $\underline{c}\notin\mathrm{Orb}(\underline{a},G)$ this sequence $g_{n}$ can be chosen from a real one parameter subgroup $H<G$. However, since $G$ is reductive, compactness of $G_{\mathbb{R}}$ implies that $G$ is $\mathbb{R}$-anisotropic i.e. contains no non-trivial $\mathbb{R}$-split tori $\mathbb{G}_{k}<G$ (see for example [borel]). Thus $H$ is trivial and therefore $\underline{c}\in\mathrm{Orb}(\underline{a},G)$. Repeating the argument for $\underline{b}$ we finally get the desired result:
$\exists g\in G:\ g\cdot\underline{a}=\underline{b}$
By Theorem˜1.1 this implies that $\underline{a}$ and $\underline{b}$ are of the same $\emptyset$-type in $\mathcal{F}$ and since all functions interpreted by $E$ are $G$-invariant, $\underline{a}$ and $\underline{b}$ are also of the same $\emptyset$-type in $(\mathcal{F},E)$. ∎

### v2

@@ 1.1 Introduction
In contrast, much less is known about quantifier elimination for noncommutative algebras. Even for natural and well-studied structures such as the classical quaternions $\mathbb{H}(k)$ or octonions $\mathbb{O}(k)$, for $k$ real closed, the model-theoretic behavior is considerably more subtle. In particular, it was recently shown in [savi] that their first-order theories in the pure ring language $\mathcal{F}$ do not admit quantifier elimination, and the problem of identifying suitable language extensions was left open.
@@
As applications, we prove quantifier elimination for the classical quaternions and octonions over real closed fields in a natural language expansion by conjugation $\bar{\ }$ and order $\leq$. More generally, we obtain criteria for finite-dimensional algebras $\mathcal{A}$ over real closed and algebraically closed fields, connecting properties of their invariant algebras $k[\mathcal{A}^{m}]^{G}$ for $G<\text{Aut}(\mathcal{A})$, to possible candidates of language extensions of $\mathcal{F}$, such that quantifier elimination holds.
@@ 1.4 Linear Algebraic Groups and Invariant Theory
Let $k$ be an algebraically closed field and $G<\mathrm{GL}_{n}(k)$. We say that $G$ is a linear algebraic group (over $k$) if it is Zariski closed in $\mathrm{GL}_{n}(k)$, i.e. if there exists an ideal $I$ of polynomial functions in the matrix entries such that
@@
The next theorem, in the form we use it, reduces to an application of the Stone-Weierstrass theorem.
@@ 2 Results
The $(\mathcal{F},\leq,\bar{\ })$-structure $\mathbb{H}(k)$ admits quantifier elimination.
@@
We apply Theorem˜2.1 to the $(\mathcal{F},\leq,\bar{\ })$-structure $\mathbb{H}(k)$ and the $(\mathcal{F},\leq)$-structure $k$, which admits quantifier elimination.
@@
each coordinate of $\iota^{(m)}(f_{m}(y_{1},...,y_{m}))$ is a simple $(\mathcal{F},\bar{\ })$-term. Therefore
@@
is a quantifier-free $(\mathcal{F},\leq,\bar{\ })$-formula defining $f_{m}^{-1}(D)$.
@@
$\mathrm{Aut}(\mathbb{H}(k))_{(\mathcal{F},\leq,\bar{\ })}=\mathrm{Aut}(\mathbb{H}(k))_{\mathcal{F}}=SO_{3}(k).$
@@
The $(\mathcal{F},\leq,\bar{\ })$-structure $\mathbb{O}(k)$ admits quantifier elimination.
@@
For both quaternions and octonions the symbol $\bar{\ }$ was needed to be able to express the invariant in our extended language. Note that $\bar{\ }$, $\operatorname{Re}$, $\operatorname{Im}$, and $N$ are interdefinable, so the precise choice of the additional symbol can be handled flexibly. On the other hand it was essential for our proof to extend $\mathcal{F}$ by $\leq$ and interpret it as the partial order on the center. This had to be done to ensure condition (ii) of Theorem˜2.1 and indeed, the symbol $\bar{\ }$ allows for defining a copy of $k$ in the $(\mathcal{F},\bar{\ })$-structure $\mathbb{H}(k)$ or $\mathbb{O}(k)$ without quantifiers. Thus formulas like
@@
Then, since $G$ is a compact subgroup of $\mathrm{Aut}(\mathcal{A})<\mathrm{GL}_{n}(k)$ it is a Lie group. Thus by Theorem˜1.5 $k[\mathcal{A}^{m}]^{G}$ separates the orbits of the diagonal action $G\curvearrowright\mathcal{A}^{m}$. Since $(s_{i})$ generate $k[\mathcal{A}^{m}]^{G}$ we get the following relation:
@@
The conditions for the language extension $E$ Proposition˜2.1 seem somewhat far fetched. However, note that so called first fundamental theorems (see [sym]) give explicit descriptions of generators for many invariant algebras of $k$-algebras, and their corresponding automorphism groups (or subgroups thereof) by means of tensor forms. These are usually stated for the algebraically closed case, however, it should be possible to often use Zariski denseness arguments, as we did in the proof of Lemma˜2.2, to infer back to the real closed case.
What about algebras $\mathcal{A}$ over an algebraically closed field (in the following always denoted $k$)? Compactness of the acting group was essential in the real closed case. Comparing with Theorem˜1.2, we see that for a general reductive linear algebraic group $G$, $k[V]^{G}$ separates only orbit closures, which need not separate $\emptyset$-types. A natural remedy for this would again be to require $G$ to be compact. In contrast to the real case ,however, compactness for reductive linear algebraic groups $G$ over $k$, only holds in the case of $G$ finite, restricting us quite a bit.
@@
First note that by reductiveness of $G$ (see below) Theorem˜1.3 all $S_{m}$ are finite and so define
@@
Since $k$ is algebraically closed, it admits quantifier elimination. Conditions (i) and (ii) of Theorem˜2.1 are verified exactly as in Proposition˜2.1.
For (iii), note that since a finite group is discrete it is by definition reductive and also obviously has closed orbits. Hence by Theorem˜1.2, $k[\mathcal{A}^{m}]^{G}$ separates $G$-orbits. As before,
@@
For an associative $k$-algebra $\mathcal{A}$, by an application of Wedderburn’s Theorem, $\text{Aut}(\mathcal{A})$ is only finite in the case $\mathcal{A}\cong\mathbb{C}^{n}$. However finite subgroups of $\text{Aut}(\mathcal{A})$ exist and their invariant algebras, can be studied ([compu, fin]). For non-associative algebras, $\text{Aut}(\mathcal{A})$ can even itself be finite (and non-trivial). For example in [low], all semisimple, two dimensional $k$-algebras have been classified and their corresponding automorphism groups (some of which are finite) and invariant algebras calculated explicitly.
For infinite reductive linear algebraic groups, the situation becomes much more hopeless. Conditions on specific points $x$ to give closed orbits exist (see for example [KN, birkes]), but these are hard to apply for the whole space simultaneously. For example if $G$ is defined over $\mathbb{R}$ with a regular representation on $V$ over $\mathbb{R}$ and $G_{\mathbb{R}}$ compact it is easy to imply by a result in [birkes] that the orbits of the real points in $V_{\mathbb{R}}$ under $G$ are closed, however, this still fails to transfer to all points in $V$.
In fact it follows immediately by the Hilbert-Mumford stability criterion [GIT], that if $G$ acts nontrivially on some complex vector space $V$ and $G$ contains a torus $\mathbb{C}^{\times}$, there will always be some $0\neq v\in V$ such that $0\in\overline{\text{Orb}(x,G)}$.